OpenStudy (anonymous):

How do i rearrange this equation for x 2x + 18/2 = 24/2

OpenStudy (anonymous):

Is the whole expression divided by 2?

OpenStudy (anonymous):

On the left side?

OpenStudy (anonymous):

2x+18/2=12 (Divide 24 by 2) 2x+18=24 (Multiply by 2 to get rid of the denominator) 2x=6 (Subtract 18) x=3 (Divide by three) or 2x+9=12 2x=3 x=3/2 or 1.5 I'm guessing the whole thing is divided by 2 (so it should be the first answer).